# Distinguishing Neutron Star vs. Low-Mass Black Hole Binaries with Late Inspiral & Postmerger Gravitational Waves — Sensitivity to Transmuted Black Holes and Non-Annihilating Dark Matter

### v2

@@ 4.2 Exclusion Sensitivity for DM-Nucleon Interaction
The TBH-TBH mergers could be observable in GW detectors as LMBH mergers. The absence of such detections thus places stringent constraints on the DM-nucleon scattering cross-section $\sigma_{\chi n}$[28, 47]. However, the constraints in Ref. [28] were obtained under the assumption that no LMBHs would be actually detected. The method presented in this paper can be used to assess the exclusion sensitivity to the DM parameter space without making the above assumption. Interpreting the LMBHs as TBHs (Eq. 4.6), we translate our limits on $f_{\rm LMBH}$ into bounds on the DM mass and its scattering cross section with nucleons, as shown in Fig. 8. For this purpose, we replace $f\times R_{\rm CBC}$ in Eq. 4.4 with the TBH-TBH merger rate $R_{\rm TBH}$ (Eq. 4.6), and determine the region in the $(m_{\chi}-\sigma_{\chi n})$ plane where the statistical significance $Z$ is $\geq 1.28$.
@@
In Fig. 8 we shows projected exclusion on the DM parameter space. For this analysis, we integrated up to a maximum luminosity distance of $D_{L}=450\,\mathrm{Mpc}$ to compute the total number of signal and background events. The resulting bounds thus reflect the contribution from all the nine bins up to 450 Mpc. The exclusion sensitivity of Fig.  8, for 10 years observation with LIGO A+, is slightly stronger than the projected upper limits in Fig. 1 of Ref. [28]. This is simply because the $\langle\rm VT\rangle$ in this case is approximately twice the exposure used in the forecasted limit in Ref. [28]. Evidently, ET can achieve stronger exclusions with 1 year of data. Although here we focused on bosonic DM, the same technique can be applied to other scenarios, including fermionic DM or bosonic DM that forms a Bose Einstein condensate (BEC) [28].
@@ 5 Discussions & Conclusion
In this study, we have systematically investigated the potential degeneracy between BNS mergers and LMBH mergers. As a first step, we considered systems with identical component masses and spin configurations for both BNS and LMBH binaries, focusing on the $_{\odot}-1.35M1.35\,M_{\odot}$ case. While the LMBH waveform can be obtained directly from standard general relativistic binary black hole models, generating accurate BNS waveforms is considerably more challenging due to the involvement of complex microphysical processes, most notably the dependence on the NS EoS. To capture this effect, we extracted BNS waveforms from the publicly available CoRe database, carefully selecting eight different EoSs ranging from the stiffest to the softest cases, and analyzed their impact on the waveform morphology, particularly in the late inspiral and postmerger phases.
@@
After obtaining the two sets of waveforms, we analyzed the mismatch between them. We found that the primary source of mismatch arises in the high-frequency regime, where strong matter effects become significant due to the increasing influence of gravity in the late inspiral and postmerger phases. This behavior is consistent with the expectation that the microphysical properties of NS matter, governed by the EoS, manifest predominantly in the strong-field regime. We divide the waveforms into two distinct phases; the inspiral and the postmerger. Through this separation, we demonstrate that the fit between the BNS and LMBH waveforms deteriorates significantly in the postmerger phase compared to the inspiral. This is expected, as the postmerger dynamics are strongly influenced by matter effects and the tidal interactions that depend sensitively on the NS EoS. These EoS-dependent effects result in diverse outcomes for BNS merger remnants, which can be short-lived or long-lived hypermassive NS, or may undergo a prompt collapse into a BH, leading to characteristic differences in the emitted GW signal. We compute the Bayesian evidence in favor of LMBH models over BNS models across a variety of NS EoSs, focusing on how this evidence evolves as a function of luminosity distance. Our analysis highlights the critical role of detector sensitivity and matter effects at different distances, which in turn influences the ability to distinguish between LMBH and BNS mergers. This variation in Bayesian evidence with distance has direct implications for key astrophysical measurements, most notably, the estimation of the BNS merger rate. Misidentification of LMBH mergers as BNS events at large distances can lead to an overestimation of the true BNS rate, thereby impacting theoretical modeling, and population studies.
@@
We then computed expected exclusions on the maximal fraction of LMBH mergers contributing to the population of CBCs. We perform a thorough statistical analysis incorporating both genuine LMBH merger signals and background contributions arising from the misclassification of BNS mergers, particularly in regimes where waveform degeneracies occur due to limited detector sensitivity. We further consider the scenario in which these LMBHs originate from DM capture-induced transmutation of BNSs. In this context, the limits on $f_{\rm LMBH}$ can be translated into constraints on the DM parameter space, specifically the DM mass $m_{\chi}$ and its scattering cross section with nucleons $\sigma_{\chi n}$. These exclusions are similarly based on rejecting the LMBH+BNS hypothesis in favor of the BNS-only hypothesis, i.e. in the absence of observed LMBH signals. We intentionally exclude the case of mixed LMBH-NS binaries in this study, as one of our motivations was LMBHs formed via dark matter capture-induced transmutation. Given that the binary neutron stars are typically close enough to reside in the same ambient DM environment, it is expected that both components undergo transmutation, resulting in symmetric LMBH-LMBH systems.
@@
To summarize, we present a numerical-relativity–based framework to distinguish BNSs from LMBHs and constrain exotic formation channels, including DM-induced scenarios such as TBH formation. Unlike prior studies relying on inspiral-only waveforms, our analysis spans both inspiral and postmerger regimes and quantifies waveform degeneracies using Bayesian evidence and Fitting Factors across detector sensitivities and distances. Prima facie the inspiral phase provides a stronger discrimination between BNSs and LMBHs than the corresponding postmerger phase. However, this may not remain true when one includes more parameters (e.g., spins). The Fitting Factors for the waveforms in the postmerger phase may be less subject to parameter degeneracies, as the difference comes from a distinct feature. This has implications for dark matter constraints and BNS rate estimates, and highlights the need for improved high-frequency modeling [125] for probing NS EoSs and exotic physics [126].
Note Added
After our work was completed, we became aware of the preprint arXiv:2507.07895[127]. The above preprint explores tidal effects in the inspiral phase of GW190425-like systems as a means to distinguish NSs from BHs. In contrast, our work presents the first systematic study of the postmerger regime, incorporating a broad set of EoSs and leveraging the information from the second peak beyond the inspiral. While the above preprint advocates for and uses a Savage-Dickey ratio to quantify distinguishability, appropriate in the case of nested hypotheses, we use the more generally applicable Bayes Factor. More broadly, while these two studies have a similar aim, they differ significantly in focus, scope, and detail. Our work has been presented previously, e.g., at PPC 2024 in IIT Hyderabad, at GW BSM 2025 in ICTS-TIFR Bengaluru, and at NDM 2025 in IHP Paris.
@@ Appendix A Appendices
A.1 Data Extraction from CoRe Database
@@
The CoRe [80] database website provides a section named ‘GW-DB’ where one can find the instructions about accessing the datasets required. One can also find information on the EoSs and a description of the data structure. The data for each waveform is stored in a HDF5 [128] file, for example named as BAMXXXX.h5. Dismantling the file with h5py [129] module of python, one can find the data structure of each waveform. The file is written in three groups:
@@
A.2 Exposure of the GW Detectors
@@
Here we present the exposure for both LIGO A+ and the Einstein Telescope (Fig. 11). The $y$-axis shows the volume-time exposure, defined as the product of the sensitive volume and observation time,
@@
Here in Fig. 11, we assume 1 year of data acquisition for each detector. The $x$-axis corresponds to redshift. For reference, a redshift of $z=0.1$ approximately corresponds to a luminosity distance ($D_{L}$) of $\sim 450\,\mathrm{Mpc}$.
A.3 Projected Exclusion Significance
Figure 12: Exclusion significance $Z$ as a function of the LMBH merger fraction $f_{\rm LMBH}$. As $f_{\rm LMBH}$ increases, the number of LMBH-induced signal events rises relative to the background, leading to a stronger statistical preference for the signal-plus-background hypothesis over the background-only hypothesis. The significance grows accordingly, allowing higher values of $f_{\rm LMBH}$ to be excluded with increased confidence for the non-observation of signal. The EoS SLy has been assumed for the BNS mergers to compute the Fitting Factor and Bayes Factor that inform expected rates.
The total number of events in the $i^{\rm th}$ bin, from a Poisson process is given by,
$P(n_{i})=\frac{(\mu S_{i}+B_{i})^{n_{i}}e^{-(\mu S_{i}+B_{i})}}{n_{i}!}\,$, (A.3)
where $S_{i}$ & $B_{i}$ in our case can be calculated using Eq. 4.3. The log-likelihood ratio of the signal hypothesis ($\mu=1$) over background-only hypothesis ($\mu=0$) is given by,
$=-2\sum_{i}\left(B_{i}\,{\rm ln}\frac{S_{i}+B_{i}}{B_{i}}-S_{i}\right)\,.$ (A.4)
We define the median expected exclusion significance $Z$, i.e. the statistical measure for rejecting the signal-plus-background hypothesis in favor of the background-only hypothesis (i.e. no signal is observed), as
${Z}=\sqrt{q}=\sqrt{2\sum_{i}\left(B_{i}\,{\rm ln}\frac{B_{i}}{S_{i}+B_{i}}+S_{i}\right)}\,.$ (A.5)
A threshold of $Z\geq 1.28$[122, 123] corresponds to a 90% confidence level, meaning that if the observed data yield $Z\geq 1.28$ (see Fig. 12), we can exclude the signal+background model at 90% confidence.

### v3

@@ 3.2 Exclusion Sensitivity for DM-Nucleon Interaction
TBH-TBH mergers could be observable in GW detectors as BLMBH mergers. The absence of such detections thus places stringent constraints on the DM-nucleon scattering cross section $\sigma_{\chi n}$[16, 35]. However, the constraints in Ref. [16] were obtained under the assumption that no BLMBHs would be actually detected. In the language of this paper, this is akin to assuming that no BNS mergers would be misclassified as BLMBH mergers. The method presented in this paper can be used to assess the exclusion sensitivity to the DM parameter space without the above assumption. Interpreting the LMBHs as TBHs (Eq. 3.9), we translate our limits on $f_{\rm BLMBH}$ into bounds on the DM mass and its scattering cross section with nucleons, as shown in Fig. 6. For this purpose, we replace $f\times R_{\rm CBC}$ in Eq. 3.7 with the TBH-TBH merger rate $R_{\rm TBH}$ (Eq. 3.9), and determine the region in the $(m_{\chi}-\sigma_{\chi n})$ plane where the statistical significance $Z$ is $\geq 1.28$.
@@
In Fig. 6 we show projected exclusion on the DM parameter space. For this analysis, we integrated up to a maximum luminosity distance of $D_{L}=450\,\mathrm{Mpc}$ to compute the total number of signal and background events. The resulting bounds thus reflect the contribution from all the nine bins up to 450 Mpc. The exclusion sensitivity of Fig.  6, for 10 years observation with LIGO A+, is slightly stronger than the projected upper limits shown in Fig. 1 of Ref. [16]. This is simply because the exposure $\langle\rm VT\rangle$ in this case is approximately twice the exposure used in the forecasted limit in Ref. [16]. ET can achieve comparable (or stronger) exclusions with 1 year of data. More broadly though, for the observation horizon considered, misclassification does not lead to a drastic degradation of the previously projected limits.
@@ 4 Discussions & Conclusion
In this study, we have systematically investigated the potential degeneracy between BNS mergers and BLMBH mergers. As a first step, we considered systems with identical component masses and spin configurations for both BNS and BLMBH binaries, focusing on the $_{\odot}-1.35M1.35\,M_{\odot}$ case. While the BLMBH waveform can be obtained directly from standard general relativistic BBH computations, generating accurate BNS waveforms is considerably more challenging due to the involvement of complex microphysical processes, most notably the dependence on the NS EoS. To capture this effect, we extracted BNS waveforms from the publicly available CoRe database, carefully selecting eight different EoSs ranging from the stiffest to the softest cases, and analyzed their impact on the late inspiral and postmerger phase waveforms.
@@
After obtaining the two sets of waveforms, we analyzed the mismatch between them. We found that the dominant mismatch arises in the high-frequency regime, as matter effects become significant in higher frequency signals from the late inspiral and postmerger phases. This behavior is consistent with the expectation that the microphysical properties of NS matter, governed by the EoS, manifest in the GW signal predominantly in the strong-field regime. We divide the waveforms into two distinct phases: the late inspiral and the postmerger. Through this separation, we demonstrate that the fit between the BNS and LMBH waveforms deteriorates significantly in the postmerger phase compared to the inspiral. The EoS-dependent effects lead to characteristic differences in the emitted GW signal. We compute the Bayesian evidence in favor of BNS models over BLMBH models, across a variety of NS EoSs. Our analysis highlights the critical role of detector’s sensitivity across low and high frequencies, which influences the ability to distinguish between LMBH and BNS mergers and which part of the signal gives more information. While natively the strongest differences lie in the postmerger phase (and detectors like NEMO can leverage it), the proposed detectors CE and ET are more sensitive at lower frequencies and the inspiral signal provides greater sensitivity therein. However, this conclusion may be altered by degeneracies with other parameters such as spins and eccentricities. The variation in Bayesian evidence with distance has direct implications for key astrophysical measurements, most notably, the estimation of the BNS merger rate. Misidentification of LMBH mergers as BNS events at large distances can lead to an overestimation of the true BNS rate, thereby impacting theoretical modeling, and population studies.
@@
We then computed expected exclusions on the maximal contribution of BLMBH mergers to the observed CBC rate. We perform a thorough statistical analysis incorporating BLMBH merger signals and background contributions arising from the misclassification of BNS mergers, particularly in regimes where waveform degeneracies occur due to limited detector sensitivity. We further consider the scenario in which these BLMBHs originate from DM capture-induced transmutation of BNSs. In this context, the limits on $f_{\rm BLMBH}$ can be translated into constraints on the DM parameter space, specifically the DM mass $m_{\chi}$ and its scattering cross section with nucleons $\sigma_{\chi n}$. These exclusions are similarly based on rejecting the BLMBH+BNS hypothesis in favor of the BNS-only hypothesis, i.e. the absence of a true BLMBH component in the low-mass CBC sample. We intentionally exclude the case of mixed BLMBH-NS binaries in this study, as one of our motivations was BLMBHs formed via dark matter capture-induced transmutation. Given that the binary neutron stars are typically close enough to reside in the same ambient DM environment, it is expected that both components undergo transmutation, resulting in symmetric LMBH-LMBH systems.
@@
To summarize, we have studied the possibility to distinguish BNSs from BLMBHs, and constrain exotic LMBH formation channels, including DM-induced scenarios such as TBH formation. Unlike prior studies relying on inspiral-only waveforms, our analysis spans both inspiral and postmerger regimes and quantifies the distinguishability using Bayesian evidence. This has implications for dark matter constraints and BNS rate estimates, and highlights the need for improved high-frequency modeling [111] for probing NS EoSs and a variety of exotic physics scenarios [112].
@@ Appendix A Appendices
In this appendix we briefly review definitions, data sources, and a number of auxiliary quantities needed to reproduce the analysis and results shown in the main text.
A.1 BLMBH Waveforms from PyCBC
[…]
We generate the LMBH-LMBH merger waveform by using the publicly available software PyCBC [68]. We employ its time domain and frequency domain module named IMRPhenomD [66, 67], which is a phenomenological model of GW signals from the inspiral-merger-ringdown of non-precessing (aligned spin) black hole binaries. These waveforms cover the entire inspiral, merger and ringdown regime, where the strong gravity effects are calibrated with numerical relativity (NR) simulations, unlike the waveforms generated using perturbative expansion in $v/c$, where $v$ is the orbital velocity and $c$ is the speed of light. Fig. 7 gives the frequency domain waveforms of LMBH mergers for different component masses.
For all our BLMBH waveforms we consider two non-spinning LMBHs with equal masses, $m_{1}=m_{2}\in 1-2.5\,M_{\odot}$, for which the chirp mass is $M_{c}=\mu^{3/5}\,M^{2/5}$, where $\mu\equiv{m_{1}m_{2}}/(m_{1}+m_{2})$ is the reduced mass of the system and $M\equiv\left(m_{1}+m_{2}\right)$ is the total mass of the system. While the IMRPhenomD waveforms we use are more detailed, below we describe the basic features of BLMBH waveforms.
The ‘plus’ (+) and ‘cross’ ($\times$) polarized strain amplitudes for the dominant ($l=2$, $m=2$) mode are given by [58]
$\tilde{h}_{\times}(f)=\left(\frac{5}{24}\right)^{1/2}\frac{1}{\pi^{2/3}}\frac{1}{D_{L}}\left(GM_{c}\right)^{5/6}f^{-7/6}e^{i\Psi_{\times}(f)}\cos{\iota},$ (A.1)
where $G$ is the gravitational constant, $D_{L}$ is the luminosity distance, $M_{c}$ is the detector-frame chirp mass given by $(1+z)M_{c}|_{\rm source\,rest\,frame}$. The redshift $z$ can be assumed to be 0 for current observations so far. The inclination of the perpendicular to the orbital plane with respect to the line of sight is given as $\iota$}. If $\iota=\pi/2$, the dominant strain amplitude is reduced by a factor of 1/2 and $h_{\times}$ vanishes, resulting in a linearly polarized wave; this orientation is termed as ‘edge-on’. For $\iota=0\,\emph{\&}\,\pi$, the system’s orbital angular momentum is aligned with the line of sight and this is termed as a ‘face-on’ orientation, resulting in a circularly polarized wave. For a face-on symmetric $\left(m_{1}=m_{2}\right)$ LMBH binary at $D_{L}=$ 1 Mpc (with vanishing spins), the strain amplitude will be
$\tilde{h}(f)\approx 4.2\times 10^{-22}\,{\rm sec}\left(\frac{\rm Mpc}{D_{L}}\right)\left(\frac{M_{c}}{1.17M_{\odot}}\right)^{5/6}\left(\frac{f}{100\,\rm Hz}\right)^{-7/6}\,$, (A.2)
with the typical $f^{-7/6}$ scaling with frequency that is expected from the inspiral of two point masses.
This analytical waveform is valid up to a maximum frequency, termed as the frequency of innermost stable circular orbit (ISCO), given by
$f_{\rm ISCO}=\frac{1}{6\sqrt{6}\left(2\pi\right)}\frac{c^{3}}{GM}\approx 2.2\,{\rm kHz}\,\left(\frac{M_{\odot}}{M}\right)\,.$ (A.3)
A.2 BNS Waveforms from CoRe Database & Choice of EoSs
Figure 8: Frequency domain waveforms of $1.35\,M_{\odot}-1.35\,M_{\odot}$ BNS mergers, with non-spinning components, for the EoSs described in Table 2.
[…]
The first release of CoRe [121] comprised 367 waveforms derived from numerical simulations based on general relativity, employing 164 distinct setups with varying parameters such as total binary mass, mass ratio, initial separation, eccentricity, and stellar spins. The second release [72] included 254 distinct BNS configurations, resulting in a total of 590 numerical-relativity simulations performed at various grid resolutions. These simulations encompass a wide range of possibilities, with 18 different EoSs and a wide range of parameters. These simulations are executed with two independent mesh-based methods: BAM [122, 123] and THC [124]. BAM is better adapted to treat the NS surface, multiple orbits in the inspiral, followed by the merger and postmerger. Whereas THC is preferentially used to address the microphysics of the NS, as it contains different neutrino transport schemes, different mixing and dissipation mechanisms during the merger. For this work, we chose 6 EoSs from BAM simulations and 2 from THC simulations.
The extraction of BNS waveforms from this database is discussed in detail (with the file structure and nomenclature) in the next section. This data is in time-domain, and we perform a Fourier transform to obtain the waveform in frequency domain. A sample of BNS waveforms from CoRe is shown in Fig. 8. They roughly agree with the BLMBH waveforms at low frequencies, but exhibit qualitative differences at higher frequencies due to the matter effects. A characteristic secondary peak associated with the fundamental oscillation mode of the remnant is noticed in the postmerger spectrum of the frequency domain. [125] Depending on the EoS and the maximum mass of the NS allowed by that EoS, the BNS merger remnant may result in a hypermassive NS that subsequently collapses to a BH, or undergo a prompt collapse [82, 81]. The frequency of this second peak (and potentially more such peaks) depends on the details postmerger. However, note the second peak seen for all the EoSs, at $f\lesssim 4$ kHz, except for 2B which is in tension with NS mass-radius data [72].
$\chi_{\rm eff}=\frac{m_{1}\chi_{1}+m_{2}\chi_{2}}{M}\,$, (A.4)
$\tilde{\Lambda}=\frac{16}{13}\left[\frac{\left(m_{1}+12m_{2}\right)m_{1}^{4}\Lambda_{1}}{M^{5}}+\frac{\left(m_{2}+12m_{1}\right)m_{2}^{4}\Lambda_{2}}{M^{5}}\right]\,$, (A.5)
$h_{+}-ih_{\times}=D_{L}^{-1}\sum_{l=2}^{\infty}\sum_{m=-l}^{l}h_{lm}(t)_{\,-2}Y_{lm}(\iota,\phi)\,$, (A.6)
where $_{-2}Y_{lm}$ are the s=2s=-2 spin-weighted spherical harmonics [58]. The inclination and orbital phase, $\iota$ and $\phi$, respectively, give the inclination of the perpendicular to the orbital plane with respect to the line of sight, and the orbital phase of the binary at a particular time. We will only be interested in the $(2,2)$ mode, as for BLMBHs.
$P\left(\rho,\epsilon\right)=P_{\rm pwp}\left(\rho\right)+\left(\gamma_{\rm th}-1\right)\rho\left(\epsilon-\epsilon_{\rm pwp}\right)\,$, (A.7)
where $P_{\rm pwp}(\rho)$ is a given piecewise polytropic EoS model [127]. The specific parameters employed for the piecewise polytropic EoS are available on the CoRe website [65] and they mimic the zero temperature EoS models described in [127]. Fig. 1 of Ref. [72] shows the $m_{\rm NS}-R_{\rm NS}$ diagram and the $\Lambda-m_{\rm NS}$ diagram of these EoS models, which gives an overview of the current parameter space of EoS models. For a given mass, $M=1.4M_{\odot}$, the EoS 2H furnishes the largest radius $R^{\rm TOV}$ $\approx$ 15.21 km, and EoS 2B the smallest $\approx$ 9.75 km [72, 128]. Evidently 2H is the stiffest (high pressure for a given density) and 2B is the softest (lower pressure for a given density) EoS in this database.
Data Extraction from CoRe Database
@@ A.2 BNS Waveforms from CoRe Database & Choice of EoSs
The CoRe [65] database website provides a section named ‘GW-DB’ where one can find the instructions about accessing the datasets required. One can also find information on the EoSs and a description of the data structure. The data for each waveform is stored in an HDF5 [129] file, for example named as BAMXXXX.h5. Dismantling the file with h5py [130], one can find the data structure of each waveform. The file is written in three groups:
@@
A.3 GW Detectors & their Exposure
We have used four detectors in this study whose sensitivity curve in the frequencies of our concern is shown in Fig.1.
LIGO A+:
The upgraded version of Advanced LIGO (2.5-generation detector)[61] has the lowest sensitivity at both high and low frequencies, among the detectors we consider, limiting its ability to probe deep inspiral or postmerger dynamics compared to next-generation observatories.
NEMO:
NEMO (2.5-generation detector) is expected to operate at cryogenic temperatures and employs shorter, stiffer suspensions to reduce thermal vibrations, especially suspension thermal noise, a dominant source at high frequencies [62]. Other key improvements include the reduction of internal thermal noise from the mirror substrates and coatings, as well as minimizing shot noise, which arises from quantum fluctuations in the laser light [62]. These combined features make NEMO particularly well-suited for capturing the detailed structure of postmerger signals, enabling deeper insight into the NS EoS and other extreme matter phenomena.
Cosmic Explorer:
CE ($3^{\rm rd}$-generation detector) is a proposed ground-based GW detector designed with 40 km arm lengths, an order of magnitude larger than LIGO’s 4 km arms, offering vastly improved strain sensitivity across a broad frequency range, particularly in the low-frequency band (510005-1000 Hz), which is crucial for observing the full inspiral phase of CBC with high SNR [63, 135].
@@ A.3 GW Detectors & their Exposure
Einstein Telescope:
ET ($3^{\rm rd}$-generation detector) is a proposed underground GW observatory featuring a unique triangular configuration with 10 km arms, designed to achieve unprecedented sensitivity across a broad frequency range, particularly below 10 Hz, surpassing LIGO A+ and NEMO by mitigating low-frequency seismic, suspension and thermal noise through underground installation and cryogenic technologies [64, 136]. ET and CE are particularly powerful for studying early inspiral physics, precision cosmology, and multi-band GW astronomy in synergy with space-based detectors.
Here we present the exposure for LIGO A+ and the Einstein Telescope (Fig. 11) that we have used for computing the event rates. The $y$-axis shows the volume-time exposure, defined as the product of the sensitive volume and observation time,
@@
Here in Fig. 11, we assume 1 year of data acquisition for each detector. The $x$-axis corresponds to redshift. For reference, a redshift of $z=0.1$ approximately corresponds to a luminosity distance ($D_{L}$) of $\sim 450\,\mathrm{Mpc}$. The function $C_{\Theta}$ encodes the detection efficiency depending on mass, distance, and SNR of the source [90, 92]. $D_{L}(z)=(1+z)D_{c}(z)$ is the luminosity distance at redshift $z$, and $\rho_{0}$ and $r_{0}$ denote the SNR threshold and characteristic distance reach of a particular detector, respectively. For our analysis we have considered $\rho_{0}=8$, and two cases, $r_{0}=$ 80 Mpc for LIGO A+ and 1591 Mpc for ET, respectively [92].